OpenStudy (anonymous):

3 sqrt 13 + 7 sqrt 13 sqrt 90 - sqrt 10 y / sqrt 7 + sqrt 3

OpenStudy (anonymous):

\[3\sqrt{13}+7\sqrt{13}=10\sqrt{13}\] just as 3 apples and 7 apples is 10 apples

OpenStudy (anonymous):

Simplify and subtract.\[\sqrt {90} - \sqrt {10}=3 \sqrt {10}- \sqrt {10}=2 \sqrt {10}\]

OpenStudy (anonymous):

The way you have the third problem written\[\frac{y}{\sqrt{7}}+\sqrt{3}=\frac{\sqrt{7}y}{7}+\sqrt{3}\]I suppose you could continue to manipulate this, but it would still be a mess.